1/1

Mèng Hàorán,

The glow on western mountains

tradukita de N. N. 81

The glow on western mountains quickly sets, The moon is climbing over the eastern lake. My hair loose, I enjoy the evening cool, I lie in peace before the open window. The wind spreads lotus scent all through the air, The sound of dripping bamboo dew is clear. Although I’d like to fetch my qin and play, To my regret, there is no-one to hear. So touched by this, I think of my old friend, Throughout the night, I’m troubled by my dreams.

Traduko de la ĉina poemo “Shān guāng” de Mèng Hàorán (∗689 – †740) en la Anglan de N. N. 81.

Arg-995-2010 (2014-04-20 00:35:06)

Tiu ĉi laŭsenca angligo troviĝas en http: // www. chinese-poems. com/ m6. html .

http://www.poezio.net/version?poem-id=995&version-id=,2010